OpenStudy (anonymous):

The acceleration due to gravity on earth is 9.8m/s^2, while on jupiterit is 23.1 m/s^2 Suppose an object is being from an innitaial height of 100 meters from the surface of each planet. Find the time it takes to reach the ground on both planets round to the nearest tenth.

OpenStudy (anonymous):

List down the information you have : initial velocity : 0 for both : acceleration :9.8 on earth, 23.1 on jupiter : displacement : 100 : time, which you are supposed to find Use one of the four kinematic equations and substitute the values v=u+at v^2=u^2+2as s=ut + 1/2at^2 s=(u+v)t/2